An equilateral triangle ABC of dimensions (3x_2y), (x+6y), 2(2x-3y). Find the perimeter of the triangle
Answer by Boreal(15235)

Perimeter is sum of
3x-2y
x+6y
4x-6y, multiplying by 2
8x-2y
the sides are equal
so 3x-2y=x+6y
2x=8y
x=4y
check with the third
x+6y=4x-6y
3x=12y
x=4y
Therefore 8x-2y above is 32y-2y=30y
or y=(1/4)x
8x-2y=30x/4 or 7.5 x
Both those express the perimeter.